how do you get the answer to jen and heather went to lunch when the bill came jen discovered that she had only 8.00 heather had enough money to pay the other part 3/5 of the bill
Answer by Maths68(1474)


Let
Total Bill = $x

Given
Jen paid = $8
Heather paid = x(3/5)=$3x/5

Total Bill=Jen Paid + Heather Paid
x=8+3x/5
Multiply by 5 both sides
5x=40+3x
5x-3x=40
2x=40
2x/2=40/2
x=20

3/5 of 20 = (3/5)(20)= 3*4 = 12
Bill was for $20 Jen paid $8 and remaining $12 paid by Heather.